Given that co = R50 million, Ī = R230 million, G = R300 million, c = ¾, T = R100 million and YF = R3 000 million the equilibrium level of output and income is ____ and in order to reach the full-employment level of income, government spending must increase by
1
Expert's answer
2018-03-08T09:35:06-0500
Given that Co = R50 million, Ī = R230 million, G = R300 million, c = ¾, T = R100 million and YF = R3 000 million the equilibrium level of output and income is 2020 and in order to reach the full-employment level of income, government spending must increase by 245 million.
C = C0 + c*(Y - T),
Y = C + I + G,
Y = 50 + 0.75*(Y - 100) + 230 + 300,
0.25Y = 505,
Y = 2020 million.
Multiplier m = 1/(1 - c) = 1/(1 - 0.75) = 4, so to increase Y by 980 million government should increase G by 980/4 = 245 million.